# Explicit estimates for the logarithmic derivative and the reciprocal of the Riemann zeta-function

### v1

@@ Abstract
In this article, we improve on two explicit bounds of order $\log t$ for $\sigma$ close to $1$: $|\zeta^{\prime}(\sigma+it)/\zeta(\sigma+it)|$ and $|1/\zeta(\sigma+it)|$. Our method provides marked improvements in the constants, especially with regards to the latter. These savings come from replacing a lemma of Borel and Carathédory, with a Jensen-type one of Heath-Brown’s.
@@ 1 Introduction
In the study of the Riemann zeta function $\zeta(s)$ and related functions, one often requires upper bounds for the logarithmic derivative of $\zeta(s)$ and the reciprocal of $\zeta(s)$. For example, these come up in the Perron formula in the process of estimating the Chebyshev function $\psi(x)=\sum_{n\leq x}\Lambda(n)$, where $\Lambda(n)$ is the von Mangoldt function, or $M(x)=\sum_{n\leq x}\mu(n)$, where $\mu(n)$ is the Möbius function. The reader is referred to [25], [3], [8], [24, (3.14)] for more information on the topic.
Trudgian in [25] gives an explicit version of results in [24, §3] originally due to Landau, obtaining bounds of order $O(\log t)$. This was then extended in [13] and [6] to bounds of order $O(\log t/\log\log t)$. However, such estimates tend to have large constants, often rendering them unsuitable for practical use. It is a non-trivial task to reduce these constants. This is due to the fact that the factor of $\zeta(s)$ in the denominator means that such estimates can only be valid in a zero-free region, and the constants rapidly increase in size as we approach the boundary of the region, i.e., $\zeta(s)\to 0$.
The method in [24, §3] first estimates $|\zeta^{\prime}(s)/\zeta(s)|$, and uses the result to derive a bound for $|1/\zeta(s)|$. It relies on two applications of the Borel–Carathédory theorem. The first application serves to obtain a bound on ${\textup{Re}}(\zeta^{\prime}(s)/\zeta(s))$, which is then used in the second to arrive at the desired bound on $|\zeta^{\prime}(s)/\zeta(s)|$. As each application slightly worsens the final estimate, we were able to improve Trudgian’s results by cutting out the first application, and instead estimating ${\textup{Re}}(\zeta^{\prime}(s)/\zeta(s))$ using a lemma of Heath-Brown [12, Lemma 3.2].
While the improvements to bounds on $|\zeta^{\prime}(s)/\zeta(s)|$ given in this paper might seem minor compared to [25], we take the opportunity to point out an error in that paper. In the quantity $A$ of [25, Lemma 4], there should be a factor of $1/c$ instead of $1/2c$. This leads to the overall constants being better than they should be. For instance, in Table $2$ of said paper, which shows bounds of the form $|\zeta^{\prime}(s)/\zeta(s)|\leq R_{1}\log t$ for $\sigma\geq 1-1/W\log t$, one should have $(W,R_{1})=(6,473)$ and $(12,54.3)$, instead of $(6,382.58)$ and $(12,40.14)$, respectively. In this article, we have recovered the erroneous bounds and even slightly improved upon them.
Furthermore, we directly derive a bound on $|1/\zeta(s)|$ from ${\textup{Re}}(\zeta^{\prime}(s)/\zeta(s))$ rather than $|\zeta^{\prime}(s)/\zeta(s)|$, eliminating the need for the Borel–Carathédory theorem in this case, leading to a substantial improvement over [25]. In addition, when $\sigma=1$, the second assertion of Corollary 4 is significantly better than [2, Proposition A.2].
@@ 2.1 Properties of the Riemann zeta function
where $Y(W_{0},10^{2})=4.339$ is admissible, and other values of $Y(W_{0},t_{0})$ with corresponding $t_{0}$ are presented in Table 5, and values of $Y(W,5)$ with corresponding $W$ are presented in Table 6.
@@ 3 Some preliminaries
[13, Theorem 4][Borel–Carathéodory] Let $s_{0}$ be a complex number, and let $R$ be a positive number possibly dependent on $s_{0}$. Suppose that the function $f(s)$ is analytic in a region containing the disc $|s-s_{0}|\leq R$. Let $M$ denote the maximum of $|\textup{Re}\,f(s)|$ on the boundary $|s-s_{0}|=R$. Then, for any $r\in(0,R)$ and any $s$ such that $|s-s_{0}|\leq r$,
@@
Seeking to apply Theorem 1 with $f(s)=-\zeta(s)^{\prime}/\zeta(s)$, we obtain a bound on $M$ using the following lemma due to Heath-Brown [12, Lemma 3.2]. The lemma is related to the well-known one of Jensen’s, and also to one of Carleman’s (see [23, §3.71]).
@@
[12, Lemma 3.2] Let $f(z)$ be holomorphic for $|z-a|\leq R$, and non-vanishing both at $z=a$ and on the circle $|z-a|=R$. Let $\rho_{j}=a+r_{j}\exp(i\theta_{j})$ be the zeros of $f(z)$ in the disc, where $\rho_{j}$ has multiplicity $n_{j}$. Then
@@
Next we require some uniform bounds for $\zeta(s)$. We slightly improve [26, Corollary 1], which is proven using the Phragmén–Lindelöf principle. Note that we also use only two lines, $\sigma=1/2$ and $\sigma=1+\delta_{r}$, omitting the need for a bound at $\sigma=1$. This is because we intend to find the maximum bound over $\sigma\in[1/2,1+\delta_{r}]$, which we show occurs at $\sigma=1/2$, as our bound is decreasing in $\sigma$. Additionally, we found that further splitting the interval complicates (8) and raises the threshold of $t$ for the lemma to be valid.
@@
The proof is identical to that of [26, Lemma 2.1], except that we are more precise in estimating (7) by retaining the dependence on $\sigma$ in the exponents of $1+a_{1}$ and $1+a_{0}$. ∎
@@
We now refer to Lemma 3, taking the maximum of (7) at $\sigma=1/2$. It has been shown in [27, p. 15] that $Q_{0}=1.31$ is permissible in (6) together with
@@
In the remainder of this section, we present some estimates concerning the Riemann zeta function, including some for $\sigma>1$, which are better than the trivial ones. We would like to point out that Lemma 5 and Corollary 8, which can both be viewed as an extension of Corollary 7 inside the critical strip, are not actually required for the purposes of this paper. However, we include them here as the proofs are similar to that of Corollary 7, and at the same time we provide a brief discussion on their usefulness. Lemma 5 is an improvement on [15, Proposition 5.3.2], which proves a similar result.
@@
where $\varphi^{*}(u)$ is the periodic function obtained by extending the polynomial $\varphi(u)=u^{2}-u+1/6$ on the interval $[0,1]$. Note that (10) holds for $\sigma>0$ by analytic continuation (see [24, Theorem 3.5]).
@@
This bound can be used throughout (10). The first term in (10) can also be bounded using the estimate in [10, Lemma 2.1], so for $N\geq 2$,
@@
The proof follows the argument of [2, Proposition A.1], originally due to Backlund [1], and extends their result to $t<500$. We omit the proof since it is almost identical to that of Lemma 5. ∎
@@
Corollary 8 can be used in conjunction with other estimates for $\zeta(s)$ in the critical strip, such as Ford’s Theorem 1 in [9] or Trudgian’s Lemma 2.1 [26]. Ford’s bound, while asymptotically better, is only sharper than bounds of the form $\zeta(s)\ll\log t$ for very large $t$. For $\sigma\in[1-1/(W_{0}\log t),1]$ and $3\leq t\leq 10^{10^{5}}$, Corollary 8 is an improvement to [9, Theorem 1]. This makes it useful for results over smaller $t$.
@@
whereas degrading a Richert-type bound like Ford’s [9, Theorem 1] would give
@@
and a bound obtained via the Phragmén–Lindelöf principle like (2.4) in [26, Lem. 2.1] using estimates on the $1/2$-line (see Hiary, Patel, and Yang [14]) and estimates on the $1$-line (see Backlund [1]) would yield
@@
[20, Lemma 5.4] Let $\gamma$ denote Euler’s constant. For $\sigma>1$ we have
@@
It is likely that $\phi(\sigma,k)$ is always positive for any $\sigma>1$, although proving that would require a different approach, perhaps one akin to [20, Lemma 5.3]. Our method in Lemmas 7 and 8, while simpler, is restricted by bounds on $|\gamma_{n}|$ which loses some information about sign changes. It is worth noting that bounds sharper than (17) are available, but are more complicated to work with.
@@
Finally, we have a lower bound on $|\zeta(s)|$, for $\sigma>1$, which is better than the trivial bound $|\zeta(s)|\geq\zeta(2\sigma)/\zeta(\sigma)$ over certain ranges of $\sigma$ (for more information, see [17]). This is the $\log t$ analogue of [6, Lemma 8].
@@
Since $\sigma>1$, we use the classical non-negativity argument involving the trigonometric polynomial $2(1+\cos\theta)^{2}=3+4\cos\theta+\cos 2\theta$ (see [24, § 3.3] and [17]). That is, we use $\zeta^{3}(\sigma)|\zeta^{4}(\sigma+it)\zeta(\sigma+2it)|\geq 1$, which implies
@@
The two factors on the right-hand side can be bounded with Lemmas 6 and 7, respectively, with the aid of the bound
@@ 4.1 Assuming a partial Riemann hypothesis
Now to bound $|\zeta^{\prime}(s_{0})/\zeta(s_{0})|$, we can use the trivial bound for $\sigma>1$ [11, Lemma 70.1], which gives us
@@ 4.3 Combining both cases $t\leq H$ and $t\geq H$
Moving from a bound on the logarithmic derivative of $\zeta$ to one on the reciprocal of $\zeta$ is done in the usual way (see [24, Theorem 3.11]). It revolves around estimating the right-hand side of
@@
The estimation of the first term on the right-hand side can be improved using the trigonometric polynomial (see [2, Proposition A.2], [17]), while in the past the above integral has been estimated by bounding the real part of the logarithmic derivative by a uniform bound $|\zeta^{\prime}(s)/\zeta(s)|$. Here some savings can be made by splitting up the interval $[\sigma,1+\delta_{1}]$ and using the sharpest bounds for $|\zeta^{\prime}(s)/\zeta(s)|$ in the appropriate range, rather than one uniform bound for the entire interval (see [6, Lemma 10]).
However, we show that by using directly a bound on the real part instead, through Lemma 2, we are able to do much better. As before, the proof will be split into two cases, $t\leq H$ and $t\geq H$.

### v2

@@ Abstract
In this article, we improve on two explicit bounds of order $\log t$ for $\sigma$ close to $1$: $|\zeta^{\prime}(\sigma+it)/\zeta(\sigma+it)|$ and $|1/\zeta(\sigma+it)|$. Our method provides marked improvements in the constants, especially with regards to the latter. These savings come from replacing a lemma of Borel and Carathédory, with a Jensen-type one of Heath-Brown’s. Using an argument involving the trigonometric polynomial, we additionally provide a slight asymptotic improvement: $1/\zeta(\sigma+it)\ll(\log t)^{11/12}$.
@@ 1 Introduction
In the study of the Riemann zeta function $\zeta(s)$ and related functions, one often requires upper bounds for the logarithmic derivative of $\zeta(s)$ and the reciprocal of $\zeta(s)$. For example, these come up in the Perron formula in the process of estimating the Chebyshev function $\psi(x)=\sum_{n\leq x}\Lambda(n)$, where $\Lambda(n)$ is the von Mangoldt function, or $M(x)=\sum_{n\leq x}\mu(n)$, where $\mu(n)$ is the Möbius function. The reader is referred to [26], [4], [9], [25, (3.14)] for more information on the topic.
Trudgian in [26] gives an explicit version of results in [25, §3] originally due to Landau, obtaining bounds of order $O(\log t)$. This was then extended in [14] and [7] to bounds of order $O(\log t/\log\log t)$. However, such estimates tend to have large constants, often rendering them unsuitable for practical use. It is a non-trivial task to reduce these constants. This is due to the fact that the factor of $\zeta(s)$ in the denominator means that such estimates can only be valid in a zero-free region, and the constants rapidly increase in size as we approach the boundary of the region, i.e., $\zeta(s)\to 0$.
The method in [25, §3] first estimates $|\zeta^{\prime}(s)/\zeta(s)|$, and uses the result to derive a bound for $|1/\zeta(s)|$. It relies on two applications of the Borel–Carathédory theorem. The first application serves to obtain a bound on ${\textup{Re}}(\zeta^{\prime}(s)/\zeta(s))$, which is then used in the second to arrive at the desired bound on $|\zeta^{\prime}(s)/\zeta(s)|$. As each application slightly worsens the final estimate, we were able to improve Trudgian’s results by cutting out the first application, and instead estimating ${\textup{Re}}(\zeta^{\prime}(s)/\zeta(s))$ using a lemma of Heath-Brown [13, Lemma 3.2].
While the improvements to bounds on $|\zeta^{\prime}(s)/\zeta(s)|$ given in this paper might seem minor compared to [26], we take the opportunity to point out an error in that paper. In the quantity $A$ of [26, Lemma 4], there should be a factor of $1/c$ instead of $1/2c$. This leads to the overall constants being better than they should be. For instance, in Table $2$ of said paper, which shows bounds of the form $|\zeta^{\prime}(s)/\zeta(s)|\leq R_{1}\log t$ for $\sigma\geq 1-1/W\log t$, one should have $(W,R_{1})=(6,473)$ and $(12,54.3)$, instead of $(6,382.58)$ and $(12,40.14)$, respectively. In this article, we have recovered the erroneous bounds and even slightly improved upon them.
Furthermore, we directly derive a bound on $|1/\zeta(s)|$ from ${\textup{Re}}(\zeta^{\prime}(s)/\zeta(s))$ rather than $|\zeta^{\prime}(s)/\zeta(s)|$, eliminating the need for the Borel–Carathédory theorem in this case, leading to a substantial improvement over [26]. In addition, when $\sigma=1$, the second assertion of Corollary 4 is significantly better than [3, Proposition A.2].
@@ 2.1 Properties of the Riemann zeta function
where $Y(W_{0},10^{3})=3.632$ is admissible, and other values of $Y(W_{0},t_{0})$ with corresponding $t_{0}$ are presented in Table 5, and values of $Y(W,5)$ with corresponding $W$ are presented in Table 6.
@@
Corollary 5.
Let $W\geq W_{0}$. In a region
$\sigma\geq 1-\frac{1}{W\log t}$
where $\zeta(\sigma+it)\neq 0$, we have for $t\geq t_{0}$,
$\left|\frac{1}{\zeta(\sigma+it)}\right|\leq Y^{\prime}(W,t_{0})(\log t)^{11/12},$
where $Y^{\prime}(W_{0},10^{3})=9.842$ is admissible, and other values of $Y^{\prime}(W,t_{0})$ with corresponding $W$ and $t_{0}$ are presented in Table 7.
In addition, we have for $\sigma=1$ and $t\geq 2$,
$\left|\frac{1}{\zeta(1+it)}\right|\leq 7.475(\log t)^{11/12}.$
@@ 3 Some preliminaries
[14, Theorem 4][Borel–Carathéodory] Let $s_{0}$ be a complex number, and let $R$ be a positive number possibly dependent on $s_{0}$. Suppose that the function $f(s)$ is analytic in a region containing the disc $|s-s_{0}|\leq R$. Let $M$ denote the maximum of $|\textup{Re}\,f(s)|$ on the boundary $|s-s_{0}|=R$. Then, for any $r\in(0,R)$ and any $s$ such that $|s-s_{0}|\leq r$,
@@
Seeking to apply Theorem 1 with $f(s)=-\zeta(s)^{\prime}/\zeta(s)$, we obtain a bound on $M$ using the following lemma due to Heath-Brown [13, Lemma 3.2]. The lemma is related to the well-known one of Jensen’s, and also to one of Carleman’s (see [24, §3.71]).
@@
[13, Lemma 3.2] Let $f(z)$ be holomorphic for $|z-a|\leq R$, and non-vanishing both at $z=a$ and on the circle $|z-a|=R$. Let $\rho_{j}=a+r_{j}\exp(i\theta_{j})$ be the zeros of $f(z)$ in the disc, where $\rho_{j}$ has multiplicity $n_{j}$. Then
@@
Next we require some uniform bounds for $\zeta(s)$. Lemma 4 slightly improves [27, Corollary 1], which is proven using the Phragmén–Lindelöf Principle (Lemma 3).
Lemma 3.
[28, Lemma 3] Let $a,b,Q$ be real numbers such that $b>a$ and $Q+a>1$. Let $f(s)$ be a holomorphic function on the strip $a\leq\textup{Re}(s)\leq b$ such that
$|f(s)|<C\exp(e^{k|t|})$
for some $C>0$ and $0<k<\tfrac{\pi}{b-a}$. Suppose further that there are $A,B,\alpha_{1},\alpha_{2},\beta_{1},\beta_{2}\geq 0$ such that $\alpha_{1}\geq\beta_{1}$ and
$|f(s)|\leq\begin{cases}A|Q+s|^{\alpha_{1}}\log^{\alpha_{2}}|Q+s|\quad\text{for}\textup{Re}(s)=a;\\ B|Q+s|^{\beta_{1}}\log^{\beta_{2}}|Q+s|\quad\text{for }\textup{Re}(s)=b.\end{cases}$
Then for $a\leq\textup{Re}(s)\leq b$, one has
$|f(s)|\leq\Big{(}{A|Q+s|^{\alpha_{1}}\log^{\alpha_{2}}|Q+s|}\Big{)}^{\tfrac{b-\textup{Re}(s)}{b-a}}\Big{(}{B|Q+s|^{\beta_{1}}\log^{\beta_{2}}|Q+s|}\Big{)}^{\tfrac{\textup{Re}(s)-a}{b-a}}.$
@@
The proof is analogous to that of [27, Lemma 2.1], except that we are more precise in estimating (7) by retaining the dependence on $\sigma$ in the exponents of $1+a_{1}$ and $1+a_{0}$.
Note that different to the proof of [27, Lemma 2.1], we use only two lines, $\sigma=1/2$ and $\sigma=1+\delta_{r}$, omitting the need for a bound at $\sigma=1$. This is because we intend to find the maximum bound over $\sigma\in[1/2,1+\delta_{r}]$, which we show occurs at $\sigma=1/2$, as our bound is decreasing (by the condition (8)) in $\sigma$. Additionally, we found that further splitting the interval complicates (8) and raises the threshold of $t$ for the lemma to be valid. ∎
@@
We now refer to Lemma 4, taking the maximum of (7) at $\sigma=1/2$. It has been shown in [29, p. 15] that $Q_{0}=1.31$ is permissible in (6) together with
@@
Lemma 5.
For every $|t|\geq 3$, we have
$|\zeta(1+it)|\leq 58.096(\log|t|)^{2/3}.$
Proof.
We use same argument in the first paragraph of [28, p. 2], which modifies Ford’s proof in [10]. The only difference is that we now apply the argument to [2, §5], which itself is an improvement on Ford’s result. This proves the lemma. ∎
Lemma 6.
For $t\geq t_{0}\geq 3$ and $\sigma\geq 1$, we have
$|\zeta(\sigma+it)|\leq C_{3}(\log t)^{2/3},$
with
$C_{3}=58.096\sqrt{1+9t_{0}^{-2}}\left(1+\frac{\log\sqrt{1+9t_{0}^{-2}}}{\log t_{0}}\right).$ (10)
Proof.
The proof is similar to that of [14, Lemma 7]. First, to show that the result holds for $1\leq\sigma\leq 2$, we use the Phragmén–Lindelöf Principle, i.e., Lemma 3, on the holomorphic function
$f(s)=(s-1)\zeta(s).$
To this end, on the $1$-line we have
$|f(1+it)|=|t|\,|\zeta(1+it)|\leq 58.096\,|2+it|(\log|2+it|)^{2/3}.$
This inequality is verified numerically for $|t|<3$, and is a consequence of Lemma 5 for $|t|\geq 3$.
Also, on the line $s=2+it$,
$|f(2+it)|\leq|1+it|\,\zeta(2)<58.096|3+it|(\log|3+it|)^{2/3},$
for all real $t$. Now apply Lemma 3 with the following parameter values: $a=1$, $b=2$, $Q=1$, $\alpha_{1}=\beta_{1}=1$, $\alpha_{2}=\beta_{2}=2/3$, and $A=B=58.096$. We obtain
$|f(s)|\leq 58.096|1+s|(\log|1+s|)^{2/3},$ (11)
for $1\leq\sigma\leq 2$. Moreover, for the same range of $\sigma$, we have the following estimates:
$\begin{split}\frac{|1+s|}{|s-1|}&\leq\frac{\sqrt{9+t^{2}}}{t}\leq\sqrt{1+9t_{0}^{-2}},\\ \log|1+s|&\leq\log\sqrt{9+t^{2}}\leq\log t+\log\sqrt{1+9t_{0}^{-2}}.\end{split}$
Thus for $1\leq\sigma\leq 2$ and $t\geq t_{0}$, the inequality (11) implies that
$|\zeta(\sigma+it)|$ $\leq 58.096\sqrt{1+9t_{0}^{-2}}\left(1+\frac{\log\sqrt{1+9t_{0}^{-2}}}{\log t_{0}}\right)\log^{2/3}t<C_{3}\log^{2/3}t,$
as desired, where $C_{3}$ is defined in (10).
Finally, we address the case where $\sigma\geq 2$. Here since $\sigma>1$, we apply the trivial bound in (48). Then for $\sigma\geq 2$, it is easy to verify that
$|\zeta(\sigma+it)|\leq 2\leq C_{3}(\log t)^{2/3}$
for all $t\geq t_{0}$. ∎
@@
In the remainder of this section, we present some estimates concerning the Riemann zeta function, including some for $\sigma>1$, which are better than the trivial ones. We would like to point out that Lemma 8 and Corollary 9, which can both be viewed as an extension of Corollary 8 inside the critical strip, are not actually required for the purposes of this paper. However, we include them here as the proofs are similar to that of Corollary 8, and at the same time we provide a brief discussion on their usefulness. Lemma 8 is an improvement on [16, Proposition 5.3.2], which proves a similar result.
@@
where $\varphi^{*}(u)$ is the periodic function obtained by extending the polynomial $\varphi(u)=u^{2}-u+1/6$ on the interval $[0,1]$. Note that (12) holds for $\sigma>0$ by analytic continuation (see [25, Theorem 3.5]).
@@
This bound can be used throughout (12). The first term in (12) can also be bounded using the estimate in [11, Lemma 2.1], so for $N\geq 2$,
@@
The proof follows the argument of [3, Proposition A.1], originally due to Backlund [1], and extends their result to $t<500$. We omit the proof since it is almost identical to that of Lemma 8. ∎
@@
Corollary 9 can be used in conjunction with other estimates for $\zeta(s)$ in the critical strip, such as Ford’s Theorem 1 in [10] or Trudgian’s Lemma 2.1 [27]. Ford’s bound, while asymptotically better, is only sharper than bounds of the form $\zeta(s)\ll\log t$ for very large $t$. For $\sigma\in[1-1/(W_{0}\log t),1]$ and $3\leq t\leq 10^{10^{5}}$, Corollary 9 is an improvement to [10, Theorem 1]. This makes it useful for results over smaller $t$.
@@
whereas degrading a Richert-type bound like Ford’s [10, Theorem 1] would give
@@
and a bound obtained via the Phragmén–Lindelöf principle like (2.4) in [27, Lem. 2.1] using estimates on the $1/2$-line (see Hiary, Patel, and Yang [15]) and estimates on the $1$-line (see Backlund [1]) would yield
@@
We also remark that although Lemma 6 is asymptotically superior to Corollary 8, one would expect that the former beats the latter only for a sufficiently large $t$; approximately $t\geq\exp(58^{3})\geq 10^{84736}$, say. This can be improved by lowering the constant in Lemma 5, and correspondingly $C_{3}$.
@@
[21, Lemma 5.4] Let $\gamma$ denote Euler’s constant. For $\sigma>1$ we have
@@
It is likely that $\phi(\sigma,k)$ is always positive for any $\sigma>1$, although proving that would require a different approach, perhaps one akin to [21, Lemma 5.3]. Our method in Lemmas 10 and 11, while simpler, is restricted by bounds on $|\gamma_{n}|$ which loses some information about sign changes. It is worth noting that bounds sharper than (19) are available, but are more complicated to work with.
@@
Finally, we have a lower bound on $|\zeta(s)|$, for $\sigma>1$, which is better than the trivial bound $|\zeta(s)|\geq\zeta(2\sigma)/\zeta(\sigma)$ over certain ranges of $\sigma$ (for more information, see [18]). This is the $\log t$ analogue of [7, Lemma 8].
@@
Since $\sigma>1$, we use the classical non-negativity argument involving the trigonometric polynomial $2(1+\cos\theta)^{2}=3+4\cos\theta+\cos 2\theta$ (see [25, § 3.3] and [18]). That is, we use $\zeta^{3}(\sigma)|\zeta^{4}(\sigma+it)\zeta(\sigma+2it)|\geq 1$, which implies
@@
The two factors on the right-hand side can be bounded with Lemmas 9 and 8, respectively, with the aid of the bound
@@ 4.1 Assuming a partial Riemann hypothesis
Now to bound $|\zeta^{\prime}(s_{0})/\zeta(s_{0})|$, we can use the trivial bound for $\sigma>1$ [12, Lemma 70.1], which gives us
@@ 4.3 Combining both cases $t\leq H$ and $t\geq H$
Moving from a bound on the logarithmic derivative of $\zeta$ to one on the reciprocal of $\zeta$ is done in the usual way (see [25, Theorem 3.11]). It revolves around estimating the right-hand side of
@@
The estimation of the first term on the right-hand side can be improved using the trigonometric polynomial (see [3, Proposition A.2], [18]), while in the past the above integral has been estimated by bounding the real part of the logarithmic derivative by a uniform bound $|\zeta^{\prime}(s)/\zeta(s)|$. Here some savings can be made by splitting up the interval $[\sigma,1+\delta_{1}]$ and using the sharpest bounds for $|\zeta^{\prime}(s)/\zeta(s)|$ in the appropriate range, rather than one uniform bound for the entire interval (see [7, Lemma 10]).
However, we show that by using directly a bound on the real part instead, through Lemma 2, we are able to do much better. Furthermore, through a trick utilising the trigonometric polynomial, we are able to get an slight asymptotic improvement on the bounds, at the cost of a slightly worse constant. As before, the proofs will be split into two cases, $t\leq H$ and $t\geq H$.